peter selects a month for his holiday.find the probability that will be april or december.assume that all months have an equal opportunity of being selected


1
Expert's answer
2021-03-09T08:08:36-0500
"P(April)=\dfrac{1}{12}, P(December)=\dfrac{1}{12}"

These two events  are disjoint, then the probability of either event is the sum of the probabilities of the two events:


"P((April)\cup (December))"

"=P(April)+P(December)"

"=\dfrac{1}{12}+\dfrac{1}{12}=\dfrac{1}{6}"

The probability that will be april or december is "\dfrac{1}{6}."